OpenStudy (anonymous):

Using the given zero, find one other zero of f(x). Explain the process you used to find your solution. 1 - 2i is a zero of f(x) = x4 - 2x3 + 6x2 - 2x + 5.

OpenStudy (helder_edwin):

do u know what is the conjugate of a complex number, for example 1-2i?

OpenStudy (anonymous):

nope

OpenStudy (helder_edwin):

ok. if u have a complex number \[ \large z=a+ib \] then its conjugate is \[ \large \overline{z}=a-ib \] u have a complex root z=1-2i. what would its conjugate be?

OpenStudy (anonymous):

i have no clue :(

OpenStudy (helder_edwin):

its conjugate would be 1+2i u just have to change the sign of whatever is with the \(i\)

OpenStudy (anonymous):

ohh ok so first change the sign way after

OpenStudy (helder_edwin):

now. when a polynomial has a complex root (such as 1-2i in your case) then its conjugate is also a root of the polynomial (1+2i in this case).

OpenStudy (helder_edwin):

this means that your polynomial can factor into \[ \large f(x)=(x-(1-2i))(x-(1+2i))g(x) \] ok?

OpenStudy (anonymous):

okk

OpenStudy (helder_edwin):

u just have to find g(x) in whichever way u know

OpenStudy (anonymous):

which is the best way to do it the easiest?

OpenStudy (helder_edwin):

it depends on u: either long division or synthetic division

OpenStudy (helder_edwin):

whatever u find easiest

OpenStudy (anonymous):

can show me how to do synthetic

OpenStudy (helder_edwin):

it's gonna take long to do it with the keyboard

OpenStudy (helder_edwin):

give me a second

OpenStudy (anonymous):

ok then long division if its easier

OpenStudy (anonymous):

what ever works for u

OpenStudy (helder_edwin):

i m gonna show u synthetic division just give a minute to type everything correctly

OpenStudy (anonymous):

ok would it be esier to draw

OpenStudy (helder_edwin):

no way

OpenStudy (helder_edwin):

\[ \large \begin{array}{c|ccccc} & 1 & -2 & 6 & -2 & 5\\ 1-2i & & 1-2i & -5 & 1-2i & -5\\ \hline & 1 & -1-2i & 1 & -1-2i & 0\\ 1+2i & & 1+2i & 0 & 1+2i\\ \hline & 1 & 0 & 1 & 0 \end{array} \]

OpenStudy (helder_edwin):

this means that \[ \large g(x)=x^2+1 \]

OpenStudy (helder_edwin):

got it?

OpenStudy (anonymous):

thats all?

OpenStudy (helder_edwin):

what r the roots (zeros) of g(x)? that's what u r asked for

OpenStudy (anonymous):

2

OpenStudy (helder_edwin):

no

OpenStudy (helder_edwin):

\[ \large 0=g(x)=x^2+1 \] then \[ \large x^2=-1 \]

OpenStudy (anonymous):

so -1

OpenStudy (anonymous):

or -1/2

OpenStudy (helder_edwin):

no :( \[ \large x=\pm\sqrt{-1}=\pm i \]

OpenStudy (anonymous):

just i?

OpenStudy (helder_edwin):

+i and -i it's a quadratic equation

OpenStudy (anonymous):

ohh so 

OpenStudy (helder_edwin):

\[ \large \pm\ \sqrt{-1} \]

OpenStudy (helder_edwin):

u have two solutions

OpenStudy (anonymous):

i have no clue wats the other 1

OpenStudy (anonymous):

wait i thought i only find 1 of the other zero

OpenStudy (helder_edwin):

the four zeros (roots) of the polynomial are \[ \large 1-2i\qquad\text{given} \] \[ \large 1+2i \] \[ \large i \] \[ \large -i \]

OpenStudy (anonymous):

right

OpenStudy (helder_edwin):

so u r done!!

OpenStudy (anonymous):

ohhhh

OpenStudy (helder_edwin):

did u understand everything?

OpenStudy (anonymous):

ya thanks to you ur the BEST

OpenStudy (helder_edwin):

u r welcome